OpenStudy (anonymous):

Solve for x: |x + 9| = 17 A. {8, 26} B. {8, -26} C. {-8, 26} D. {-8, -26}

jimthompson5910 (jim_thompson5910):

If |x + 9| = 17 then x + 9 = 17 or x + 9 = -17

jimthompson5910 (jim_thompson5910):

solve each equation for x to get your two answers

OpenStudy (anonymous):

thanks bud!

jimthompson5910 (jim_thompson5910):

np